# Topology of Discrete Quantum Feedback Control

### v1

@@ VII.3 QND demon
Figure 15: Eigenspectrum of a quantum channel for feedback control with strong unitary symmetry. In each figure, blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC in the infinite-size limit, and orange dots constitute the eigenspectrum under the OBC with the number of unit cells $L=40$. (a) Full eigenspectrum of the CPTP map $\tilde{\mathcal{E}}$. (b) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$, which is identical to that of $\tilde{\mathcal{E}}_{\downarrow\downarrow}$. (c) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\downarrow}$. (d) Eigenspectrum of $\tilde{\mathcal{E}}_{\downarrow\uparrow}$. The paramaters of the model are set to $J_{1}=J_{2}=1$, $V=+\infty$, $h_{z}=0.5$, and $h_{s}=1$. The duration of feedback is $\tau=2$.
@@ VII.4 $\mathbb{Z}_{2}$ demon
So far, we have discussed topological feedback controls in class AI, ${}_{-}$, and A, all of which are characterized by an integer-valued winding number in 1D. In contrast, the topological classification for class ${}^{\dagger}$ implies the existence of a $\mathbb{Z}_{2}$ topological invariant in 1D [34]. Here we present topological feedback control that is characterized by a $\mathbb{Z}_{2}$ topological invariant.
As shown in Sec. VI.3.3, a quantum channel in class ${}^{\dagger}$ requires unitary symmetry and $C$ symmetry with $\eta_{C}=-1$ of the truncated Bloch matrix. To construct feedback control that fulfills these requirements, we consider a spin-$1/2$ particle on a 1D lattice with two sublattices as in Sec. VII.3. We first perform a spin-nondisturbing projective measurement of the position and the sublattice, where the measurement operator is given by Eq. (143). For a measurement outcome $m_{1}=(j_{1},a_{1})$, we perform feedback control with Hamiltonian $\hat{H}_{m_{1}}$ that has a symmetry
@@
$\hat{U}_{m_{1},m_{2}}^{\prime}=\begin{cases}\exp[-i\pi\hat{S}^{z}]&(j_{1}>j_{2}\ \mathrm{and}\ a_{1}\neq a_{2});\\ \hat{I}&(\mathrm{otherwise}).\end{cases}$ (154)
@@
~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}=$ $U_{m_{1},m_{2}}^{\prime}\hat{M}_{j_{2},a_{2},\sigma}\hat{U}_{m_{1}}\hat{M}_{j_{1},a_{1},\sigma}$
@@
The unitary operator with the feedback Hamiltonian $H_{m_{1}}$ statisfies the following reciprocity relations because of the symmetry between $\hat{H}_{j_{1},A}$ and $\hat{H}_{j_{1},B}$:
@@
which follows from the symmetry (150) of the feedback Hamiltonian. Then, the conditions (110a)-(110c) are satisfied, and thus ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ with $\sigma\neq\sigma^{\prime}$ belongs to class ${}^{\dagger}$, while that with $\sigma=\sigma^{\prime}$ belongs to class ${}_{+}$.
The topological feedback control in class ${}^{\dagger}$ exhibits the $\mathbb{Z}_{2}$ skin effect [72]. To see this, we show the eigenspectrum of the quantum channel for the $\mathbb{Z}_{2}$ demon under the PBC and the OBC in Fig. 17. The block-diagonalized CPTP map $\tilde{\mathcal{E}}_{\uparrow\uparrow}$ does not possess $C$ symmetry with $\eta_{C}=-1$, and therefore the Kramers degeneracy of eigenvalues is not ensured. Consequently, the loop structure in the eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$ shown in Fig. 17(a) collapses near the real axis due to real-complex transitions and does not support nontrivial topology. In contrast, the block $\tilde{\mathcal{E}}_{\uparrow\downarrow}$ has $C$ symmetry with $\eta_{C}=-1$, leading to a Kramers-degenerated eigenspectrum shown in Fig. 17(b). The point-gap topology of this spectrum is characterized by a $\mathbb{Z}_{2}$ topological invariant [72]
@@
Figure 17: Eigenspectrum of a quantum channel for the $\mathbb{Z}_{2}$ demon. In each figure, blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC in the infinite-size limit, and orange dots constitute the eigenspectrum under the OBC with the number of unit cells $L=20$. (a) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$. (b) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\downarrow}$. The paramaters of the model are set to $J_{1}=0.8,J_{2}=1$ and, $\lambda=0.5$. The duration of feedback is $\tau=1$.
@@
VIII Dissipative feedback control
VIII.1 Formalism

### v2

@@ VII.3 QND demon
Figure 16: Eigenspectrum of a quantum channel for feedback control with strong unitary symmetry. In each figure, blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC in the infinite-size limit, and orange dots form the eigenspectrum under the OBC with the number of unit cells $L=40$. (a) Full eigenspectrum of the CPTP map $\tilde{\mathcal{E}}$. (b) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$, which is identical to that of $\tilde{\mathcal{E}}_{\downarrow\downarrow}$. (c) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\downarrow}$. (d) Eigenspectrum of $\tilde{\mathcal{E}}_{\downarrow\uparrow}$. The paramaters of the model are set to $J_{1}=J_{2}=1$, $V=+\infty$, $h_{z}=0.5$, and $h_{s}=1$. The duration of feedback is $\tau=2$.
@@ VII.4 $\mathbb{Z}_{2}$ demon
So far, we have discussed topological feedback controls in classes AI, AI $+$ psH-, and A, all of which are characterized by an integer-valued winding number in 1D. In contrast, the topological classification for class AII implies the existence of a $\mathbb{Z}_{2}$ topological invariant in 1D [34]. Here we present topological feedback control that is characterized by a $\mathbb{Z}_{2}$ topological invariant.
As shown in Sec. VI.3.3, a quantum channel in class AII requires unitary symmetry and $C$ symmetry with $\eta_{C}=-1$ of the truncated Bloch matrix. To construct feedback control that fulfills these requirements, we consider a spin-$1/2$ particle on a 1D lattice with two sublattices as in Sec. VII.3. We first perform a spin-nondisturbing projective measurement of the position and the sublattice, where the measurement operator is given by Eq. (144). For a measurement outcome $m_{1}=(j_{1},a_{1})$, we perform feedback control with Hamiltonian $\hat{H}_{m_{1}}$ that has a symmetry
@@
$\hat{U}_{m_{1},m_{2}}^{\prime}=\begin{cases}\exp[-i\pi\hat{S}^{z}]&\mathrm{if}\ j_{1}>j_{2}\ \mathrm{and}\ a_{1}\neq a_{2};\\ \hat{I}&\mathrm{otherwise}.\end{cases}$ (155)
@@
~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}=$ $\hat{U}_{m_{1},m_{2}}^{\prime}\hat{M}_{j_{2},a_{2},\sigma}\hat{U}_{m_{1}}\hat{M}_{j_{1},a_{1},\sigma}$
@@
The unitary operator with the feedback Hamiltonian $\hat{H}_{m_{1}}$ statisfies the following reciprocity relations because of the symmetry between $\hat{H}_{j_{1},A}$ and $\hat{H}_{j_{1},B}$:
@@
which follows from the symmetry (151) of the feedback Hamiltonian. Then, conditions (111a)-(111c) are satisfied, and thus ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ belongs to class AII for $\sigma\neq\sigma^{\prime}$ and class AI $+$ psH+ for $\sigma=\sigma^{\prime}$.
The topological feedback control in class AII exhibits the $\mathbb{Z}_{2}$ skin effect [72]. To see this, we show the eigenspectrum of the quantum channel for the $\mathbb{Z}_{2}$ demon under the PBC and the OBC in Fig. 18. The block-diagonalized CPTP map $\tilde{\mathcal{E}}_{\uparrow\uparrow}$ does not possess $C$ symmetry with $\eta_{C}=-1$, and therefore the Kramers degeneracy of eigenvalues is not ensured. Consequently, the loop structure in the eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$ shown in Fig. 18(a) collapses near the real axis due to real-complex transitions and does not support nontrivial topology. In contrast, the block $\tilde{\mathcal{E}}_{\uparrow\downarrow}$ has $C$ symmetry with $\eta_{C}=-1$, leading to the Kramers-degenerated eigenspectrum shown in Fig. 18(b). The point-gap topology of this spectrum is characterized by a $\mathbb{Z}_{2}$ topological invariant [72]
@@
Figure 18: Eigenspectrum of a quantum channel for the $\mathbb{Z}_{2}$ demon. In each figure, blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC in the infinite-size limit, and orange dots form the eigenspectrum under the OBC with the number of unit cells $L=20$. (a) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$. (b) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\downarrow}$. The paramaters of the model are set to $J_{1}=0.8,J_{2}=1$ and, $\lambda=0.5$. The duration of feedback is $\tau=1$.
@@
The distinct topology of $\mathcal{E}_{\sigma\sigma}$ and $\mathcal{E}_{\sigma\sigma^{\prime}}$ $(\sigma^{\prime}\neq\sigma)$ can also be detected through dynamics. Suppose that the initial density matrix of the system is given by $\hat{\rho}_{\mathrm{i}}=\ket{j_{0},a_{0},+}\bra{j_{0},a_{0},+}$ |, where |$\ket{j_{0},a_{0},+}=(\ket{j_{0},a_{0},\uparrow}+\ket{j_{0},a_{0},\downarrow})/\sqrt{2}$ is a site-localized state with its spin polarized along the $x$ direction. Then, the particle-number density and the spin density of the final state after $N$ feedback cycles are given by
$\mathrm{Tr}[\hat{n}_{j,a}\mathcal{E}^{N}(\hat{\rho}_{\mathrm{i}})]=$ $\frac{1}{2}(\bra{j,a,\uparrow}\mathcal{E}_{\uparrow\uparrow}^{N}(\hat{\rho}_{\mathrm{i}})\ket{j,a,\uparrow}$
$+\bra{j,a,\downarrow}\mathcal{E}_{\downarrow\downarrow}^{N}(\hat{\rho}_{\mathrm{i}})\ket{j,a,\downarrow}),$ (164a)
$\mathrm{Tr}[\hat{S}_{j,a}^{x}\mathcal{E}^{N}(\hat{\rho}_{\mathrm{i}})]=$ $\mathrm{Re}[\bra{j,a,\uparrow}\mathcal{E}_{\uparrow\downarrow}^{N}(\hat{\rho}_{\mathrm{i}})\ket{j,a,\downarrow}],$ (164b)
$\mathrm{Tr}[\hat{S}_{j,a}^{y}\mathcal{E}^{N}(\hat{\rho}_{\mathrm{i}})]=$ $\mathrm{Im}[\bra{j,a,\uparrow}\mathcal{E}_{\uparrow\downarrow}^{N}(\hat{\rho}_{\mathrm{i}})\ket{j,a,\downarrow}],$ (164c)
$\mathrm{Tr}[\hat{S}_{j,a}^{z}\mathcal{E}^{N}(\hat{\rho}_{\mathrm{i}})]=$ $\frac{1}{4}(\bra{j,a,\uparrow}\mathcal{E}_{\uparrow\uparrow}^{N}(\hat{\rho}_{\mathrm{i}})\ket{j,a,\uparrow}$
$-\bra{j,a,\downarrow}\mathcal{E}_{\downarrow\downarrow}^{N}(\hat{\rho}_{\mathrm{i}})\ket{j,a,\downarrow}),$ (164d)
where
$\hat{n}_{j,a}=$ $\sum_{\sigma}\ket{j,a,\sigma}\bra{j,a,\sigma},$ (165a)
$\hat{S}_{j,a}^{x}=$ |$\frac{\ket{j,a,\uparrow}\bra{j,a,\downarrow}+\ket{j,a,\downarrow}\bra{j,a,\uparrow}}{2},$ (165b)
$\hat{S}_{j,a}^{y}=$ |$\frac{\ket{j,a,\uparrow}\bra{j,a,\downarrow}-\ket{j,a,\downarrow}\bra{j,a,\uparrow}}{2i},$ (165c)
$\hat{S}_{j,a}^{z}=$ |$\frac{\ket{j,a,\uparrow}\bra{j,a,\uparrow}-\ket{j,a,\downarrow}\bra{j,a,\downarrow}}{2}.$ (165d)
Thus, topology of the off-diagonal sectors $\mathcal{E}_{\sigma\sigma^{\prime}}$ $(\sigma^{\prime}\neq\sigma)$ is reflected in the dynamics of the transverse components of spin, whereas the dynamics of the number density and the longitudinal component of spin are governed by the diagonal sectors $\mathcal{E}_{\sigma\sigma}$.
In Fig. 19, we plot the number density and the transeverse spin components of a particle after $N$ repeated applications of the feedback control with the $\mathbb{Z}_{2}$ demon. The numerical results clearly show that the transverse spin components are localized at the edges due to the skin effect, whereas the number density is delocalized over the system. Thus, the $\mathbb{Z}_{2}$ skin effect of the topological feedback control manifests itself only in the off-diagonal components of the density matrix. Note that the transverse spin components are not conserved due to the additional feedback in Eq. (155) and vanish in the steady state of the quantum channel [see Fig. 19 (g) and (h)]. Therefore, the nontrivial topology of the off-diagonal sectors appear only in the transient dynamics under repeated applications of the feedback control.
Figure 19: Dynamical signatures of nontrivial topology of a $\mathbb{Z}_{2}$ demon. Physical quantities after $N$ feedback cycles are plotted. (a) [(b)] Number density in sublattice $A$ ($B$) for $N=40$. (c) [(d)] Density of the $x$ component of spin in sublattice $A$ ($B$) for $N=30$. (e) [(f)] Density of the $y$ component of spin in sublattice $A$ ($B$) for $N=15$. (g) [(h)] Dynamics of the total $x$ ($y$) component of spin. The initial state is given by $\hat{\rho}_{\mathrm{i}}=\ket{j_{0},A,+}\bra{j_{0},A,+}$ | with $j_{0}=L/2$ (see text). The parameters of the model are the same as in Fig. 18.
VII.5 Chern demon
In the previous subsections, we have focused on the models of quantum feedback control that are characterized by point-gap topology. In this subsection, we present an example of topological Maxwell’s demon that exhibits nontrivial line-gap topology. Our main idea is to simulate the Haldane model [101] of a Chern insulator by using feedback control. We consider a spin-$1/2$ particle on a honeycomb lattice and perform a projective measurement of the position of the particle so that the measurement does not disturb the spin state as in Sec. VII.3 and VII.4. The measurement operator is thus given by
$\hat{M}_{\vec{j},a}=\ket{\vec{j},a,\uparrow}\bra{\vec{j},a,\uparrow}+\ket{\vec{j},a,\downarrow}\bra{\vec{j},a,\downarrow},$ (166)
where $a=A,B$ denotes a sublattice of the honeycomb lattice, and we record a measurement outcome $m_{1}=(\vec{j}_{1},a_{1})$. After the measurement, we let the system undergo unitary time evolution during time $\tau$ according to the tight-binding Hamiltonian
$\hat{H}=$ $J_{1}\sum_{\sigma=\uparrow,\downarrow}\sum_{\langle\vec{i},\vec{j}\rangle}(\ket{\vec{i},A,\sigma}\bra{\vec{j},B,\sigma}+\mathrm{H.c.})$
$+J_{2}\sum_{\sigma=\uparrow,\downarrow}\sum_{\langle\langle\vec{i},\vec{j}\rangle\rangle}\sum_{a=A,B}(\ket{\vec{i},a,\sigma}\bra{\vec{j},a,\sigma}+\mathrm{H.c.}),$ (167)
where $J_{1}\in\mathbb{R}$ and $J_{2}\in\mathbb{R}$ are the hopping amplitudes for nearest-neighbor and next-nearest-neighbor sites, respectively. Note that this Hamiltonian $\hat{H}$ does not depend on a measurement outcome. Then, we perform the second projective measurement described by the measurement operator (166) to get an outcome $m_{2}=(\vec{j}_{2},a_{2})$. The set of measurement outcomes $\{m_{1},m_{2}\}$ informs us of the initial and final positions of the particle travelling under the Hamiltonian $\hat{H}$. Using this information of measurement outcomes, we apply feedback control with the unitary operator $\hat{U}_{m_{1},m_{2}}^{\prime}=\exp[-i\phi_{m_{1},m_{2}}\hat{S}^{z}]$, where $\phi_{m_{1},m_{2}}\in\mathbb{R}$ and $\hat{S}^{z}$ is the magnetization obtained by replacing $i$ with $\vec{i}$ in Eq. (145).
The quantum channel $\mathcal{E}$ that describes this quantum feedback control has the decomposition
$\mathcal{E}=\mathcal{E}_{\uparrow\uparrow}+\mathcal{E}_{\uparrow\downarrow}+\mathcal{E}_{\downarrow\uparrow}+\mathcal{E}_{\downarrow\downarrow}$ (168)
similarly to Eq. (156), and each block is given by
[…]
$=$ $\sum_{m_{1},m_{2}}p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1})\ket{\vec{j}_{2},a_{2},\sigma}\bra{\vec{j}_{1},a_{1},\sigma}\hat{\rho}\ket{\vec{j}_{1},a_{1},\sigma}\bra{\vec{j}_{2},a_{2},\sigma}$ | (169)
and
$\mathcal{E}_{\sigma\sigma^{\prime}}(\hat{\rho})=$ $\sum_{m_{1},m_{2}}e^{-i\phi_{m_{1},m_{2}}\mathrm{sgn}[\sigma]}p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1})\ket{\vec{j}_{2},a_{2},\sigma}\bra{\vec{j}_{1},a_{1},\sigma}\hat{\rho}\ket{\vec{j}_{1},a_{1},\sigma^{\prime}}\bra{\vec{j}_{2},a_{2},\sigma^{\prime}},$ (170)
where $\sigma^{\prime}\neq\sigma$, $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1}):=|\bra{\vec{j}_{2},a_{2},\sigma}e^{-i\hat{H}\tau}\ket{\vec{j}_{1},a_{1},\sigma}|^{2}$, and $\mathrm{sgn}[\sigma]$ is defined by $\mathrm{sgn}[\uparrow]=-\mathrm{sgn}[\downarrow]=+1$. Equation (170) indicates that one can imprint an arbitrary phase factor $e^{-i\phi_{m_{1},m_{2}}}$ on the transition amplitude $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1})$ by using feedback control. From a general perspective, this gives a significant advantage to realize topological phases with feedback control. In the present case, we set $\phi_{m_{1},m_{2}}$ to be nonzero if and only if $m_{1}$ and $m_{2}$ are nearest-neighbor sites, and we choose $\phi_{m_{1},m_{2}}=\pm\phi\in\mathbb{R}$, where the plus (minus) sign is taken if the hopping from $m_{1}$ to $m_{2}$ is in the counterclockwise (clockwise) direction. This choice of the phase factor resembles that in the Haldane model [101], thereby making the topologically trivial model (167) topological. Because of the phase factor, the block $\mathcal{E}_{\sigma\sigma^{\prime}}$ belongs to class A, while the block $\mathcal{E}_{\sigma\sigma}$ belongs to class AI due to the modular conjugation symmetry.
In Fig. 20, we plot the dispersion relation of the blocks $\mathcal{E}_{\uparrow\uparrow}$ and $\mathcal{E}_{\uparrow\downarrow}$ of the Chern demon constructed above. As seen from the figure, the phase $\phi$ imprinted by the feedback control opens a gap in the eigenspectrum of $\mathcal{E}_{\uparrow\downarrow}$, and gapless edge modes that connect the bulk bands emerge under the OBC. This is analogous to the physics of the Haldane model. Thus, the quantum channel is characterized by nontrivial line-gap topology of $\mathcal{E}_{\uparrow\downarrow}$, which results in the emergence of the boundary-localized edge modes.
Figure 20: Dispersion relations of the real part of eigenvalues of (a) the block $\mathcal{E}_{\uparrow\uparrow}$ and (b) the block $\mathcal{E}_{\uparrow\downarrow}$ of a quantum channel for the Chern demon. The parameters of the model are set to be $J_{1}=1,J_{2}=0.3,\tau=0.5$, and $\phi=\pi/2$. The unit of length is chosen to be the distance between nearest-neighbor sites. We employ the armchair edge in which one direction is periodic and the other direction is open with ten unit cells. The horizontal axis is the momentum along the edge. The eigenvalues of edge-localized states are colored in red in panel (b). The imaginary part of the eigenvalues is negligible.
Here, a special attention should be paid on a physical difference between a Chern insulator and the Chern demon presented in this section. A Chern insulator is characterized by line-gap topology of a Hamiltonian, which is a generator of the dynamics of a system. Since the derivative of the dispersion relation with respect to momentum gives the group velocity, the gapless spectrum of edge states leads to chiral modes, which propagate unidirectionally along the edges. In contrast, the Chern demon is defined from line-gap topology of a quantum channel, which is a time-evolution operator itself rather than its generator. Thus, the physical meaning of the gapless edge-mode spectrum of the Chern demon is different from that of a Chern insulator. In particular, the boundary-localized modes of the Chern demon do not necessarily propagate along the edges of the system. Its physical implication merits further investigation.
VIII Topological phenomena arising from complete positivity and trace preservation
Although topological feedback control exploits non-Hermitian topology of quantum channels, it is crucially different from quantum systems with non-Hermitian Hamiltonians due to the fact that a quantum channel must satisfy complete positivity and trace preservation, i.e., it is a CPTP map [47]. Here, we discuss how these properties lead to topological phenomena that are absent in conventional non-Hermitian systems.
VIII.1 Boundary-localized steady state
Complete positivity and trace preservation ensure that a quantum channel gives a map between density matrices, which leads to the existence of a steady state with unit eigenvalue [102]. Consequently, the topological Maxwell’s demons discussed in Sec. IV and VII (except for those in Sec. VII.4 and VII.5) exhibit boundary-localized steady states under the OBC. This is in stark contrast to topological phases described by non-Hermitian Hamiltonians, where the conservation of the norm of a state is not guaranteed and therefore a steady state does not exist in general. Although those boundary-localized steady states in the topological feedback control appear to be closely related to the non-Hermitian skin effect, we can construct a model that shows the non-Hermitian skin effect with no localized steady state (see Appendix F). In this sense, the emergence of boundary-localized steady states is a special feature of quantum channels and should be distinguished from the non-Hermitian skin effect. We note that a similar phenomenon is observed in classical stochastic processes where the probability conservation is satisfied unlike general non-Hermitian systems [103].
VIII.2 Breakdown of conventional bulk-boundary correspondence in non-Hermitian skin effects
The bulk-boundary correspondence in non-Hermitian systems with nontrivial point-gap topology states that a nonzero winding number of the eigenspectrum under the PBC leads to boundary-localized skin modes under the OBC whose eigenvalues generically reside inside the region delimited by the PBC spectrum [72]. However, eigenvalues of the boundary-localized steady states discussed in Sec. VIII.1 must be unity by definition and do not belong to the internal region enclosed by the PBC spectrum (see e.g., Figs. 6, 8, and 13). More strikingly, it is possible that some eigenvalues under the OBC lie outside the region enclosed by the PBC spectrum (see the case of the QND demon shown in Fig. 16). These facts indicate that the conventional bulk-boundary correspondence in non-Hermitian skin effects does not hold in topological feedback control.
The breakdown of the conventional bulk-boundary correspondence is attributed to the requirement of complete positivity and trace preservation for a quantum channel. The Hamiltonian under the OBC in a conventional non-Hermitian system is obtained by eliminating the matrix elements of the PBC Hamiltonian across the boundary of the system. However, such a straightforward implementation of the OBC for a quantum channel violates complete positivity and trace preservation. Instead, a quantum channel under the OBC should be defined as follows. Suppose that a quantum channel for feedback control under the PBC is given by
$\mathcal{E}_{\mathrm{PBC}}(\hat{\rho})=\sum_{m}e^{-i\hat{H}_{m,\mathrm{PBC}}\tau}\hat{M}_{m}\hat{\rho}\hat{M}_{m}^{\dagger}e^{i\hat{H}_{m,\mathrm{PBC}}\tau},$ (171)
where $\hat{H}_{m,\mathrm{PBC}}$ is a feedback Hamiltonian under the PBC. Then, the corresponding quantum channel under the OBC is defined as
$\mathcal{E}_{\mathrm{OBC}}(\hat{\rho})=\sum_{m}e^{-i\hat{H}_{m,\mathrm{OBC}}\tau}\hat{M}_{m}\hat{\rho}\hat{M}_{m}^{\dagger}e^{i\hat{H}_{m,\mathrm{OBC}}\tau},$ (172)
by using time evolution governed by the Hamiltonian $\hat{H}_{m,\mathrm{OBC}}$ under the OBC so that $\mathcal{E}_{\mathrm{OBC}}$ gives a CPTP map. By comparing Eqs. (171) and (172), one can see that the matrix elements of the quantum channel near the boundary are reconstructed to ensure complete positivity and trace preservation.
Since the OBC for a quantum channel is defined in a manner different from that for a Hamiltonian, the existing proofs of the bulk-boundary correspondence in non-Hermitian systems [33, 72, 73, 88] cannot simply be applied to quantum channels. The bulk-boundary correspondence in topological feedback control should be formulated so as to incorporate complete positivity and trace preservation.
VIII.3 Topological distinction between steady states and transient dynamics
The complete positivity also impacts the topological classification. As discussed in Sec. V.2, the complete positivity (or Hermiticity preservation following therefrom) imposes the modular conjugation symmetry on a quantum channel. While some symmetry sectors of a quantum channel do not necessarily satisfy the modular conjugation symmetry (as shown in Sec. VI.3), a sector that includes a steady state must satisfy it since the steady-state density matrix must be Hermitian [93]. Thus, symmetry classes allowed for a steady-state sector of a quantum channel are restricted to those three classes (AI, AI $+$ psH+, and AI $+$ psH-) that include the modular conjugation symmetry. This means that topological classification for boundary-localized steady states in Sec. VIII.1 is more restrictive than the general topological classification of non-Hermitian systems due to the complete positivity.
If transient dynamics are considered, topological phenomena that are prohibited in steady states due to complete positivity become possible. An example is given by the $\mathbb{Z}_{2}$ demon in Sec. VII.4. As shown in Fig. 18, a spin-diagonal sector $\mathcal{E}_{\uparrow\uparrow}$ that includes a steady state is topologically trivial, whereas an off-diagonal sector $\mathcal{E}_{\uparrow\downarrow}$ displays nontrivial point-gap topology owing to the absence of the modular conjugation symmetry. The nontrivial topology in the off-diagonal sector manifests itself in the transient dynamics of the transverse components of the spin density, which vanish in the steady state (see Fig. 19). The topological distinction between the steady state and transient dynamics thus enriches topological phenomena in quantum feedback control.
IX Dissipative feedback control
IX.1 Formalism